OpenStudy (tainted):

Solve the rational equation, Medal will be given 1/40 + 1/25 = 1/x Eric and Lance both play a video game, It takes Eric 40 minutes to beat the round. But it only takes Lance about 25 minutes.

OpenStudy (tainted):

\[\frac{ 1 }{ 40 }+\frac{ 1 }{ 25 } = \frac{ 1 }{ x }\]

OpenStudy (zehanz):

What is the Least Common Multiple of 40 and 25?

OpenStudy (tainted):

5

OpenStudy (zehanz):

Multiples of 40: 40, 80, 120, ... Multiples of 25: 25, 50, 75, ... LCM(40, 25)= ?

OpenStudy (zehanz):

BTW, 5 is the Greatest Common Divisor of 40 and 25...

OpenStudy (tainted):

Oops got confused for a second, It is 200 for LCM

OpenStudy (zehanz):

OK, so that is used for you new denominator: \(\dfrac{...}{400}+\dfrac{...}{400}=\dfrac{1}{x}\)

OpenStudy (zehanz):

OOPS! Make that 200 instead of 400 :(

OpenStudy (zehanz):

Giving you \(\dfrac{5}{200}+\dfrac{8}{200}=\dfrac{1}{x}\)

OpenStudy (tainted):

So\[\frac{ 40 }{ 200 } + \frac{ 25 }{ 200 } = \frac{ }{ x }\]

OpenStudy (tainted):

Oop's forgot to simplify the numerators

OpenStudy (zehanz):

No, to get from denominator 40 to 200, you multiplied it with 5, so do the same with the numerator.

OpenStudy (zehanz):

Also to get fron denomminator 25 to 200, you had to multiply with 8, so do the same with the numerator...

OpenStudy (zehanz):

That is why you get \(\dfrac{5}{200}+\dfrac{8}{200}=\dfrac{13}{200}=\dfrac{1}{x}\)

OpenStudy (tainted):

Once I get to 5/200+8/200=13/200=1/x how do I solve it

OpenStudy (anonymous):

13x=200 x=200/13

OpenStudy (zehanz):

You almost have the solution: if \(\dfrac{1}{x}=\dfrac{13}{200}\), then \(\dfrac{x}{1}=\dfrac{200}{13}\), so \(x=\dfrac{200}{13} \approx...\)